OpenStudy (anonymous):

what would a cosine function look like with an amplitude of 7 period of pie/6 and a vertical shift down of 2

OpenStudy (anonymous):

@ganeshie8

OpenStudy (goformit100):

▬Sir/Ma'am, A Warm Welcome To Open Study. Sir/Ma'am please show your Working for the question you have posted.▬

OpenStudy (anonymous):

@ganeshie8

OpenStudy (kropot72):

\[y-A \cos (Bx+C)+D\] |A| is the amplitude. The period is \[T=\frac{2 \pi}{B}\] The phase shift is \[\frac{C}{B}\] and the vertical shift is D. Can you plug the given values into the above equation?

OpenStudy (anonymous):

Sure but can you check if I am correct or not once I do

OpenStudy (kropot72):

You will need to find the value of B by solving \[\frac{2 \pi}{B}=\frac{\pi}{6}\]

OpenStudy (anonymous):

what would be B though

OpenStudy (anonymous):

like i know i have to solve but do i cross multiply and divide in this case

OpenStudy (kropot72):

You can solve in steps: 1. Multiply both sides by B. 2. Multiply both sides of the result by 6/pi.

OpenStudy (anonymous):

So for the first step it would look like this 2piB/bb?

OpenStudy (kropot72):

Actually just cross-multiplying is easier. Would you like to try that?

OpenStudy (anonymous):

Yes please

OpenStudy (anonymous):

Would the answer be 2?

OpenStudy (kropot72):

Cross-multiplying gives: \[B \times \pi=12\times \pi\] Now, divide both sides by pi to find B. Can you do that?

OpenStudy (anonymous):

B=12

OpenStudy (kropot72):

Good work! So can you now plug the values into the general equation, but omit the phase shift because no phase shift is given in the question?

OpenStudy (kropot72):

There is a typo in my general equation. It should be (omitting the phase shift this time): \[y=A \cos Bx+D\]

OpenStudy (kropot72):

A = 7 B = 12 D = -2